# Learning to crawl: benefits and limits of centralized vs distributed control

### v2

@@ III.3 Robustness
We next show that centralized architectures are considerably more robust to failure of individual suckers. In fig. 4, we focus on the best policy obtained for each architecture for the 12 suckers crawler and test its robustness to failure. To this end, we extract one or more random suckers at each integration step, and let them perform a random action (adhere or not with the same probability). All other suckers follow the optimal policy, and no extra learning is allowed. Centralized architectures are more robust to failure: their performance drops slower with the number of failing suckers fig. 4(a), with a single random failure eroding less than 10% performance vs about 20% for distributed architectures, fig. 4(b). We then asked how robust are these policies to failure of the most crucial sucker. When we make a single sucker fail consistently over time, we find that the head plays the most prominent role in all of our architectures, as its failure contributes the largest drop in performance fig. 4(c). In fact the head is always the dominant sucker when the CPG wavelength, $\lambda$, is equal or larger than the crawler length and not otherwise (appendix C). Failure of the head is particularly acute for distributed architectures (60% drop in performance for fully distributed architectures) against much less dramatic drops in centralized architectures (the most robust being the 2CC with a $<20\%$ drop).
@@
To further quantify the distinct policies we measure the frequency that each sucker experiences the four local compression/elongation states of its adjacent springs, as the policy is played over time. The four local states are represented with the symbols $\rightarrow$|$\rightarrow$ | →; $\leftarrow$|$\rightarrow$ | →; $\rightarrow$|$\leftarrow$ | ← and $\leftarrow$|$\leftarrow$ | ←. The local states are visited with frequencies that vary across the different architectures (fig. 4 (e) top row), but the pattern is relatively preserved. This suggests that the underlying physics constrains the spring dynamics considerably, and may partly overshadow differences in the control architecture. In contrast, the average action played by each sucker in its local compression/elongation state varies dramatically across different control architectures (fig. 4 (e) bottom row). In distributed architectures (left panels), suckers play an action that depends only on their local compression/elongation state, hence colors are binary (adhere or not adhere). In contrast, in centralized architecture, a sucker acts in response to the global (albeit binary) state of compression/elongation, and not only in response of its local state. Intermediate colors in the right panels signify that the local compression/elongation state is not enough to determine whether the sucker will adhere or not. Because centralized policies are dictated by global compression/elongation states, they may be more tolerant to failures of individual suckers.
@@
Figure 5: Distribution of speed for the policies learned by the 12 suckers crawler. Top left: number of distinct policies learned by the 12-sucker crawler with different architectures (blue); number of times the most popular policy was learned (red); Grey inset: detail of policies learned by the fully centralized agent; top: plateau in $G$, showing the policy $\pi_{1}$ is re-learned several times (same as bottom right panel in Fig 2b); bottom: histogram of the velocities across all 250 policies (orange) compared with the histogram of policies (blue). Bins where the blue histogram is lower than the orange histogram represent that few policies within the bin are re-learned multiple times.
@@ III.4 Redundance
Figure 6: Spring states observed when playing the best policy for the different architectures considered in the 12 suckers crawler, compared to a random deterministic policy. a) Kymographs of states experienced by each spring (x axis) during integration: each row corresponds to a different time; time increases on the y axis from top to bottom. b) Unique binary states of compression/elongation of the 11 springs experienced as the crawler adopts specific policies. Ordering of the states is chosen to guide the eye – it does not account for the time spent in each state nor the temporal order these states are encountered.
@@ III.5 Orchestrating a traveling wave
The results described so far suggest that centralized control achieves better performance at the cost of exploring a higher dimensional space of possible states and actions. Why is centralization beneficial? Our crawler is composed of individual suckers that only have the ability to sense a binary compressed/elongated state of their adjacent springs. This makes for a system of poor proprioceptors that do not have access to e.g. their spatial location nor to a clock. We find that by orchestrating the action of many poor proprioceptors, centralized control is able to ride the traveling wave generated by the CPG, despite the lack of a direct measure of space or time. To characterize this intuition, we visualize kymographs of the state of compression of the springs generated by the dynamics as a crawling policy is enacted (fig. 6(a)). Note that the state of compression of the springs is not deterministically controlled by the adhesion pattern of the suckers. We find that all optimal policies generate periodic states of compression as seen by repeating patterns representing traveling waves of compression with speed $\approx 0.19108$, perfectly matching the underlying central pattern generator phase velocity, $v_{p}/x_{0}=\omega/k\approx 0.19099$, with $\omega=0.1,\,\,k=2\pi/N_{s}$ and $N_{s}=12$. This is in contrast to random deterministic policies, which generate random states of compression of the springs and cannot reproduce the CPG wave (bottom panel of fig. 6(a)). Interestingly, for distributed architectures and especially in the hive case the bands abruptly end near the center of the crawler, signifying that when the compression wave hits springs 5 and 6 it stops: a clump of springs near the center of the crawler remains constantly elongated. In contrast, for centralized architectures the wave never stops but rather travels along the crawler all the way to the end, so that the CPG movement is effectively transferred to the crawling. Thus by orchestrating adhesion of multiple suckers at a time, centralized architectures realize a smoother traveling wave of compression. Consistently, when enacting optimal centralized policies, the dynamics generates a limited number of compression states. For the optimal fully centralized policy (1CC), the 12 springs experience only 26 states of compression, about $1\%$ of the available $2^{N_{s}-1}$ combinations of binary states associated to the $N_{s}-1$ springs. Interestingly, with distributed architectures the number of states visited using optimal policies is about 100% of the available $4(N_{s}-1)$ states (fig. 6(b), standard).
@@ IV Discussion and perspectives
Here we presented a toy model of a one dimensional crawler composed of blocks connected by springs. Each block is a rudimentary actuator representing a sucker making a single binary decision at each time: adhere vs not-adhere. The suckers are also rudimentary proprioceptors, as they measure the binary state of compression vs elongation of their adjacent springs. Positional information and speed of the individual suckers is not measured nor inferred. We assume that a central pattern generator, or CPG, periodically modifies the rest length of the springs, according to a traveling wave. Using this toy model we have demonstrated that this group of rudimentary proprioceptors/actuators can learn to interact with the substrate and exploit the wave generated by the CPG to achieve a net unidirectional motion of the crawler. Particularly, we focused on the pros and cons of distributed vs centralized forms of learning and control.
@@
Using tabular Q-learning, we found that locomotion can be learned in a pure multiagent setting, where each sucker is considered an independent agent and with no explicit communication among agents, consistent with previous literature [14]. In this setting, coupling is learned and is mediated by reward, as all suckers contribute to the dynamics that displaces the center of mass which produces reward. A purely distributed form of learning and control is cost effective in that it requires no communication and thus scales well with the number of agents and the degrees of freedom in the system.
@@
We find that introducing a certain degree of centralization, in the form of Control Centers encompassing several suckers, provides two benefits. On the one hand, the increased complexity of centralized policies can be effectively exploited to couple the actions of many suckers at a time. These long range correlations enable centralized policies to more effectively ride the wave of compression provided by the CPG, and ultimately yield faster crawling. On the other hand, as multiple suckers are activated at a time, failure of an individual sucker results in less dramatic loss of performance, making centralized policies both more effective and more robust to failure.
@@
Centralization of course comes at the cost of increased computational cost and requires two-way communication of proprioceptive input from the suckers to the control center, and of motor control from the control center to the suckers. We find that intermediate architectures where multiple control centers orchestrate the actions of a subset of suckers achieve the best of both worlds, providing virtually identical benefits of full centralization, but limiting the number of degrees of freedom to a manageable size. Ultimately, partial or full centralization overcome the limits of binary proprioception and activation by realizing smooth and continuous dynamics, despite binary input and output. Further work is needed to establish the optimal degree of centralization, emerging from the trade-offs between robustness, speed and computational cost, which will be relevant to understand the control architecture of biological crawlers and to design efficient robotic crawlers (see e.g. [22, 23]).
@@
In contrast, previous biologically inspired work focused on either a fully centralized [13] or fully distributed (hive) control [14]. In [13], a biologically plausible model of crawling is proposed which also uses rudimentary forms of proprioception and control: input is limited to the identity of the most compressed body segment and actions correspond the designation of the sole active neuron in the neural network controlling muscular contraction, while adhesion is not controlled. The authors show with a fully centralized Q–learning architecture that the central pattern generator can be learned, rather than imposed as we do here.
@@ Acknowledgments
Appendix A Materials and Methods
A.1 Equation of motion
Referring to the definitions introduced in the main text and fig. 1, the general differential equation governing the internal blocks (excluding tail and head) is given by
$\begin{split}m\ddot{x}_{n}(t)=-\zeta_{n}\dot{x}_{n}(t)+\kappa\left(x_{n+1}(t)-x_{n}(t)-l_{n}(t)\right)&\\ -\kappa\left(x_{n}(t)-x_{n-1}(t)-l_{n-1}(t)\right)\end{split}$ (7)
where $x_{n}$ is the position of the $n$th sucker, with $n=1,\dots,N_{s}-1$. For the tail, $n=0$ , and the head, $n=N_{s}$, the dynamics is given by:
$m\ddot{x}_{0}(t)$ $=-\zeta_{0}\dot{x}_{0}(t)+\kappa\left(x_{1}(t)-x_{0}(t)-l_{0}(t)\right)$ (8)
$m\ddot{x}_{N_{s}}(t)$ $=-\zeta_{0}\dot{x}_{N_{s}}(t)-\kappa\left(x_{N_{s}}(t)-x_{N_{s}-1}(t)-l_{N_{s}-1}(t)\right)$ (9)
A.2 Scaling and parametrization
Recalling the CPG equation eq. 1, $l_{n}=x_{0}+a\sin(\omega t-kn)$, being $x_{0}$ the rest position of each spring, it fixes the spatial scaling of the problem. Introducing the rescaled position of the suckers/block, $\bar{x}=x/x_{0}$, $\bar{l}=l/x_{0}$, and the constants $A=m/(\zeta_{0}\omega^{2})$ and $B=\kappa/(\zeta_{0}\omega)$, the above equations can be rewritten as:
$\begin{split}&A\ddot{\bar{x}}_{n}(t)=-\dot{\bar{x}}_{n}(t)+B\big{(}\bar{x}_{n+1}(t)+\bar{x}_{n-1}(t)-2\bar{x}_{n}(t)-\bar{l}_{n}(t)+\bar{l}_{n-1}(t)\big{)}\\ &A\ddot{\bar{x}}_{0}(t)=-\dot{\bar{x}}_{n}(t)+B\left(\bar{x}_{n+1}(t)-\bar{x}_{n}(t)-\bar{l}_{n}(t)\right)\\ &A\ddot{\bar{x}}_{N_{s}}(t)=-\dot{\bar{x}}_{n}(t)+B\left(\bar{x}_{n-1}(t)-\bar{x}_{n}(t)+\bar{l}_{n-1}(t)\right)\end{split}$ (10)
In the following, we assume the overdamped limit, that is the inertial term is negligible compared to the other terms, $A\ll B$. Therefore the differential equation representing the dynamics of the nth internal crawler in absence of adhesion simplifies to a cinematic equation:
$\dot{\bar{x}}_{n}(t)=B\left(\bar{x}_{n+1}(t)+\bar{x}_{n-1}(t)-2\bar{x}_{n}(t)-\bar{l}_{n}(t)+\bar{l}_{n-1}(t)\right)$ (11)
When the $n$th suckers is adhering to the substrate ($\zeta_{n}=\infty$) $B=0$, we simply have null instantaneous velocity, $\dot{\bar{x}}_{n}=0$. We think that the overdamped approximation is reasonable for the purpose of this work. In Appendix B.2, for validation purposes, we show the damped dynamics already approaching the overdamped limit when the ratio between inertia and friction is of order one, $m/\zeta_{0}=1$. The spatial periodicity $N$ in the CPG oscillation is in principle an arbitrary integer. For simplicity, we fix $N$ identical to the number of suckers, $N_{s}$, so that a single wavelength, $\lambda=Nx_{0}$, is contained in the crawler. Therefore, given the crawler length $L$, we have $\lambda=L$ and $x_{0}$ is scaled according to the number of suckers considered, $x_{0}=L/N_{s}$. Other parameters are fixed to $\kappa=1\,\mathrm{N}/\mathrm{m}$, $\zeta_{0}=1\,\mathrm{N}/\mathrm{m}$, $\omega=0.1\mathrm{s}^{-1}$, $a=x_{0}/4$ (the amplitude of the oscillation of the CPG), and $L=10\,\mathrm{m}$. To summarize the problem is dictated by 3 nondimensional parameters, $N_{s}=L/x_{0}$, $L/\lambda$ and $B=\kappa/(\zeta_{0}\omega)$:
[…]
Appendix B Validation
B.1 Periodic crawler
@@
Figure 9: Comparison between eq. 4.9 of [15] (eq. 4 in the main text without constant prefactors) (blue lines), and the implementation of eq. 3 in the simulator (orange dots) for the periodic crawler model. a) A more honest comparison with the perturbation regime used in Ref. [15] which accounts for the small adhesion limit, using $\epsilon=1$ in the simulator. b) With infinite adhesion in the simulator as the one used generally in this work.
@@
However, when the crawler is perturbed by the adhesion pulse of eq. 3, the perturbative analytical solution proposed by [15], we are probably bounded in the comparison by the impossibility of realizing a genuine continuous limit and an infinite tentacle. Indeed, as shown in fig. 9, the implementation of the optimal crawling strategy of eq. 3 in the periodic crawler model discussed in this appendix, does not compare well to the perturbation analysis leading to eq. 4.
However, as also shown in the Results section for the finite crawler, we have a fairly good qualitative comparison with the theoretical behavior, namely the existence of a maximum in $N$ ($=N_{s}$, the number of suckers for our parametrization), and a decay for $N\rightarrow\infty$. The position of the maximum is not exactly the one predicted by the analytical theory but similar (with $\omega=0.1$, $N_{\max}^{finite}\sim 12<\tilde{N}_{\max}=(2\pi)/\sqrt{\omega}=20$). Interestingly, as illustrated by fig. 3 in the main text, the functional shape of the dependency with $N_{s}$ is preserved even when crawling policies allow for the adhesion of multiple suckers at a time which was excluded from the analytical treatment of Ref. [15].
B.2 Damped limit
Figure 10: head position as a function of the integration time in scaled units $\bar{x}=x/x_{0}$. The damped regime collapses on the overdamped limit for values of the scaled constant $A\omega^{2}=m/\zeta_{0}<1$ in eq. 10, where the overdamped dynamics is given by eq. 11.
For validation purposes, we can drop the overdamped assumption and solve the full damped equation for the finite crawler eq. 10. In fig. 10, we show that the overdamped regime is already approached when the mass of the ”sucker block” and the friction coefficient are of the same order. Preliminary results also suggest that policies learned in the damped regime are no significantly different from the overdamped ones. Therefore, we expect that the conclusions reached in this work should still hold in the damped scenario. However, further verifications would be needed to assert this claim unequivocally.
Appendix C Effect of CPG periodicity over sucker importance

### v3

@@ III.3 Robustness
We next show that centralized architectures are considerably more robust to failure of individual suckers. In fig. 4, we focus on the best policy obtained for each architecture for the 12 suckers crawler and test its robustness to failure. To this end, we extract one or more random suckers at each integration step, and let them perform a random action (adhere or not with the same probability). All other suckers follow the optimal policy, and no extra learning is allowed. Centralized architectures are more robust to failure: their performance drop more slowly with the number of failing suckers fig. 4(a), with a single random failure eroding less than 10% performance vs about 20% for distributed architectures, fig. 4(b). We then asked how robust these policies are to failure of the most crucial sucker. When we make a single sucker fail consistently over time, we find that the head plays the most prominent role in all of our architectures, as its failure contributes the largest drop in performance fig. 4(c). In fact the head is always the dominant sucker when the CPG wavelength, $\lambda$, is equal or larger than the crawler length and not otherwise (appendix C). Failure of the head is particularly acute for distributed architectures (60% drop in performance for fully distributed architectures) against much less dramatic drops in centralized architectures (the most robust being the 2CC with a $20\%$ drop).
@@
To further quantify the distinct policies we measure the frequency that each sucker experiences the four local compression/elongation states of its adjacent springs, as the policy is played over time. The four local states are represented with the symbols $\rightarrow|\rightarrow$; $\leftarrow|\rightarrow$; $\rightarrow|\leftarrow$ and $\leftarrow|\leftarrow$. The local states are visited with frequencies that vary across the different architectures (fig. 4 (e) top row), but the pattern is relatively preserved. This suggests that the underlying physics constrains the spring dynamics considerably, and may partly overshadow differences in the control architecture. In contrast, the average action played by each sucker in its local compression/elongation state varies dramatically across different control architectures (fig. 4 (e) bottom row). In distributed architectures (left panels), suckers play an action that depends only on their local compression/elongation state, hence colors are binary (adhere or not adhere). In contrast, in centralized architecture, a sucker acts in response to the global (albeit binary) state of compression/elongation, and not only in response to its local state. Intermediate colors in the right panels signify that the local compression/elongation state is not enough to determine whether the sucker will adhere or not. Because centralized policies are dictated by global compression/elongation states, they may be more tolerant to failures of individual suckers.
@@
Figure 5: Distribution of speed for the policies learned by the 12 suckers crawler. a): number of distinct policies learned by the 12-sucker crawler with different architectures (blue); number of times the most popular policy was learned (red); b): detail of policies learned by the fully centralized agent; top: plateau in $G$, showing the optimal policy, labeled as $\pi_{1}$, is selected at several iterations of the exploration phase of the learning protocol (same as bottom right panel in fig. 2b, color change in the plateau marks policy change); bottom: histogram of the velocities across all 250 policies. Orange: number of times a crawler velocity is realized by a learned policy . Blue: number of unique policies associated to the velocity (within the bin). Bins where the blue histogram is lower than the orange histogram show identical policies which are selected multiple times in the learning protocol. c): Same as b), for the remaining learning architectures.
@@ III.4 Redundance
Figure 6: Spring states observed when playing the best policy for the different architectures considered in the 12 suckers crawler, compared to a random deterministic policy. a) Kymographs of states experienced by each spring (x axis) during integration: each row corresponds to a different time; time increases on the y axis from top to bottom. b) Unique binary states of compression/elongation of the $1$ springs experienced as the crawler adopts specific policies. States appear in an arbitrary order that does not account for the time spent in each state nor the temporal order these states are encountered (e.g. state 0 is not visited more often than state 1, nor it is visited right before state 1).
@@ III.5 Orchestrating a traveling wave
The results described so far suggest that centralized control achieves better performance at the cost of exploring a higher dimensional space of possible states and actions. Why is centralization beneficial? Our crawler is composed of individual suckers that only have the ability to sense a binary compressed/elongated state of their adjacent springs. This makes for a system of poor proprioceptors that do not have access to e.g. their spatial location nor to a clock. We find that by orchestrating the action of many poor proprioceptors, centralized control is able to ride the traveling wave generated by the CPG, despite the lack of a direct measure of space or time. To characterize this intuition, we visualize kymographs of the state of compression of the springs generated by the dynamics as a crawling policy is enacted (fig. 6(a)). Note that the state of compression of the springs is not deterministically controlled by the adhesion pattern of the suckers. We find that all optimal policies generate periodic states of compression as seen by repeating patterns representing traveling waves of compression with speed $\approx 0.19108$, perfectly matching the underlying central pattern generator phase velocity, $v_{p}/x_{0}=\omega/k\approx 0.19099$, with $\omega=0.1,\,\,k=2\pi/N_{s}$ and $N_{s}=12$. This is in contrast to random deterministic policies, which generate random states of compression of the springs and cannot reproduce the CPG wave (bottom panel of fig. 6(a)). Interestingly, for distributed architectures in the hive case the bands abruptly end near the center of the crawler, signifying that when the compression wave hits springs 5 and 6 it stops: a clump of springs near the center of the crawler remains constantly elongated. In contrast, for centralized architectures the wave never stops but rather travels along the crawler all the way to the end, so that the CPG movement is effectively transferred to the crawling. Thus by orchestrating adhesion of multiple suckers at a time, centralized architectures realize a smoother traveling wave of compression. Consistent with the observation that centralized policies generate a smooth travelling wave of compression, the dynamics does not generate all possible combinations of compression and elongation of all springs, which would be $2^{N_{s}-1}$. Instead, only a limited number of compression states is visited. More precisely, for the optimal fully centralized policy (1CC), the 12 springs experience only 26 states of compression, about $1\%$ of the available $2^{N_{s}-1}$ combinations of binary states associated to the $N_{s}-1$ springs. When the crawler enacts any other policy, the coupled dynamics generates more states, indicative of stumbling blocks in wave propagation. This is despite the fact that centralized policies respond to the global state of compression of all springs simultaneously, and are thus much more complex. Thus, somewhat counter intuitively, more complex policies with more degrees of freedom generate fewer states than simpler policies with fewer degrees of freedom.
@@ IV Discussion and perspectives
Here we presented a toy model of a one dimensional crawler composed of blocks connected by springs. Each block is a rudimentary actuator representing a sucker making a single binary decision at each time: adhere vs not-adhere. The suckers are also rudimentary proprioceptors, as they measure the binary state of compression vs elongation of their adjacent springs. Positional information and speed of the individual suckers is not measured nor inferred. We assume that a central pattern generator, or CPG, periodically modifies the rest length of the springs, according to a traveling wave. The CPG is dictated by an internal clock, independent from proprioception. This may be a useful model for the octopus, where proprioceptive feedback has been proposed to be implemented with an embodied mechanism rather than a neural process based on somatotopy [22]. Using this toy model we have demonstrated that this group of rudimentary proprioceptors/actuators can learn to interact with the substrate and exploit the wave generated by the CPG to achieve a net unidirectional motion of the crawler. Particularly, we focused on the pros and cons of distributed vs centralized forms of learning and control.
@@
Using tabular Q-learning, we found that locomotion can be learned in a pure multiagent setting, where each sucker is considered an independent agent and with no explicit communication among agents, consistent with previous literature [18]. Although the agents do not communicate explicitly, the result of their actions is indirectly coupled via the reward. Indeed, the reward is the net translocation which results from the dynamics of the system of oscillators, hence to maximize reward agents effectively learn how to act cooperatively. A purely distributed form of learning and control is cost effective in that it requires no communication and thus scales well with the number of agents and the degrees of freedom in the system.
@@
We find that introducing a certain degree of centralization, in the form of Control Centers encompassing several suckers, provides two benefits. On the one hand, the increased complexity of centralized policies can be effectively exploited to couple the actions of many suckers at a time. These long range correlations enable centralized policies to more effectively ride the wave of compression provided by the CPG, and ultimately yield faster crawling. On the other hand, as multiple suckers are activated at a time, failure of an individual sucker results in less dramatic loss of performance, making centralized policies more robust to failure.
@@
Centralization of course comes at the cost of increased computational cost and requires two-way communication of proprioceptive input from the suckers to the control center, and of motor control from the control center to the suckers. We find that intermediate architectures where multiple control centers orchestrate the actions of a subset of suckers achieve the best of both worlds, providing virtually identical benefits of full centralization, but limiting the number of degrees of freedom to a manageable size. Ultimately, partial or full centralization overcome the limits of binary proprioception and activation by realizing smooth and continuous dynamics, despite binary input and output. Further work is needed to establish the optimal degree of centralization, emerging from the trade-offs between robustness, speed and computational cost, which will be relevant to understand the control architecture of biological crawlers and to design efficient robotic crawlers (see e.g. [8, 45, 35, 2]).
@@
In contrast, previous biologically inspired work with simple models of crawling (or templates [14]) focused on either a fully centralized [36] or fully distributed (hive) control [18]. In [36], a biologically plausible model of crawling is proposed which also uses rudimentary forms of proprioception and control: input is limited to the identity of the most compressed body segment and actions correspond the designation of the sole active neuron in the neural network controlling muscular contraction, while adhesion is not controlled. The authors show with a fully centralized Q–learning architecture that the central pattern generator can be learned, rather than imposed as we do here. The benefits of distributed vs centralized control for animal locomotion have been studied in several other contexts and algorithms, and generally pertain added stability in uneven terrain, adaptation, versatility, see e.g. [59, 58, 50, 38].
@@ Acknowledgments
Appendix A Validation
A.1 Periodic crawler
@@
Figure 9: Comparison between eq. 4.9 of [51] (blue lines), and the implementation of eq. 5 in the simulator (orange dots) for the periodic crawler model. a) A more honest comparison with the perturbation regime used in Ref. [51] which accounts for the small adhesion limit, using $\epsilon=1$ in the simulator. b) With infinite adhesion in the simulator as the one used generally in this work.
@@
When the crawler is perturbed by the adhesion pulse of eq. 5, an asymptotic solution is available from ref. [51] (fig. 9). The implementation of the optimal crawling strategy (eq. 5) in the periodic crawler discussed here, has the same qualitative behavior predicted by the theory, eq. 6, with a different location and magnitude of the maximum. These discrepancies likely arise due to our approximation of the infinite crawler with the periodic crawler.
However, as also shown in the Results section for the finite crawler, we have a fairly good qualitative comparison with the theoretical behavior, namely the existence of a maximum in NN ($N_{s}$, the number of suckers for our parametrization), and a decay for $N\rightarrow\infty$. The position of the maximum is not exactly the one predicted by the analytical theory but similar (with $\omega=0.1$, $N_{\max}^{finite}\sim 12<\tilde{N}_{\max}=(2\pi)/\sqrt{\omega}=20$). Interestingly, as illustrated by fig. 3 in the main text, the functional shape of the dependency with $N_{s}$ is preserved even when crawling policies allow for the adhesion of multiple suckers at a time which was excluded from the analytical treatment of Ref. [51].
Damped limit
For validation purposes, we can drop the overdamped assumption and solve the full damped equation for the finite crawler eq. 1. The overdamped regime is already approached when inertia balances the restoring force of the spring, i.e. $m\omega^{2}/\kappa=1$. Preliminary results also suggest that policies learned in the damped regime are not significantly different from the overdamped ones. Therefore, we expect that the conclusions reached in this work should still hold in the damped scenario. However, further verifications would be needed to assert this claim unequivocally.
Appendix B Effect of CPG periodicity over sucker importance